Three different companies each purchased trucks on January 1, 2018, for $70,000. Each truck was expected to last four years or 200,000 miles. Salvage value was estimated to be $5,000. All three trucks were driven 66,000 miles in 2018, 41,000 miles in 2019, 39,000 miles in 2020, and 61,000 miles in 2021. Each of the three companies earned $59,000 of cash revenue during each of the four years. Company A uses straight-line depreciation, company B uses double-declining-balance depreciation, and company C uses units-of-production depreciation.

Answer each of the following questions. Ignore the effects of income taxes.

Problem 6-24 Part a

Required

  1. a-1. Calculate the net income for 2018? (Round "Per Unit Cost" to 3 decimal places.)

Company A:

Company B:

Company C:

  1. a-2. Which company will report the highest amount of net income for 2018?

  • Company A

  • Company B

  • Company C

  • All of the choices

b-1. Calculate the net income for 2021? (Round "Per Unit Cost" to 3 decimal places.)

Company a:

Company b:

Company c:

c-1. Calculate the book value on the December 31, 2020, balance sheet? (Round "Per Unit Cost" to 3 decimal places.)

Company a

Company b

company c

d-1. Calculate the retained earnings on the December 31, 2021, balance sheet?

Company a:

Company b:

Company c:

What can you conclude concerning the effects of the choice of different depreciation methods on the financial statements over the life of the assets? Include comments regarding net income, balance sheet amounts and cash flow.
Answer #1

Straight line Depreciation = (Original Value - Salvage Value) / Useful Life
= ($70000 - 5000) / 4 = $16250 per year

Double Declining Depreciation = Beginning Value x 2 times straight line rate

Year Beginning Value DDB @50% Ending Value
2018 $                       70,000 $           35,000 $              35,000
2019 $                       35,000 $           17,500 $              17,500
2020 $                       17,500 $             8,750 $                8,750
2021 $                         8,750 $             3,750 $                5,000

Units of Production Method = (Original Value - Salvage Value) / Total production
= ($70000 - 5000) / 207000 = $65000 / 207000 per mile

Year Miles Beginning Value Depreciation Acc Dep Ending Value
2018 66000 $           70,000 $              20,725 $              20,725 $             49,275
2019 41000 $           49,275 $              12,874 $              33,599 $             36,401
2020 39000 $           36,401 $              12,246 $              45,845 $             24,155
2021 61000 $           24,155 $              19,155 $              65,000 $                5,000
207000

a-1 Net Income for 2018 = Cash Revenue - Depreciation
Company A = $59000 - $16250 = $42750
Company B = $59000 - $35000 = $24000
Company C = $59000 - $20725 = $38275

a-2
Company A will report highest income, since depreciation is lowest

b-1 Net Income for 2018 = Cash Revenue - Depreciation
Company A = $59000 - $16250 = $42750
Company B = $59000 - $3750 = $55250
Company C = $59000 - $19155 = $39845

c-1 Book value on the December 31, 2020 = Original Value - Accumulated Depreciation
Company A = $70000 - $16250x3 = $21250
Company B = $70000 - $61250 = $8750
Company C = $70000 - $45845 = $24155

d-1 Balance of retained earnings on Dec 31, 2021 = $59000 x 4 - $65000 = $171000
Company A = $171000
Company B = $171000
Company C = $171000

Selection Depreciation method shows different incomes for each year under different methods, but mostly depreciation throughout the life of asset remains same. So it effects the timing of income only.